Brad has a cylindrical barrel with radius 10 inches and height 15 inches. He wants to fill it from a well, using a bucket in the shape of a hemisphere with a radius of 10 inches. How many trips must Brad make to the well in order to fill the barrel?
Answer by Boreal(15235)

Volume of barrel is pi*r^2*h=25*pi*15=375 pi in^3
volume of hemisphere is (2/3)*pi*r^3=(250/3)*pi in^3
divide the first by the second, and pi cancels
375/(250/3)=375*3/250=1125/250
needs five trips
=================
Check using the fact that 1 1/2 spheres or 3 hemispheres will fit into the barrel physically. That would be the minimum, but there is empty space, and that would require four full hemispheres and part of a fifth.